OpenStudy (anonymous):

1, 5, 14, 30, 55, ___, ___

OpenStudy (turingtest):

adding squares +2^2,+3^2,+4^2,+5^2 so 55+6^2=91 and91+7^2=140 so 90, 140

OpenStudy (anonymous):

what does ^ mean?

OpenStudy (turingtest):

raised to that power\[+2^2,+3^2,+4^2,...\] you will see that notation a lot, i.e. 3^3=27